OpenStudy (anonymous):

1. Find the inverse of the following function. Find the domain, range, and asymptotes of each function. Graph both functions on the same coordinate plane. f(x)= 7-e^-x

OpenStudy (anonymous):

can you guys help me please?

OpenStudy (anonymous):

\[y=7-e^{-x}\] \[x=7-e^{-y}\] solve for \(y\) \[x-7=e^{-y}\] \[\ln(x-7)=-y\] \[y=-\ln(x-7)\]

OpenStudy (anonymous):

@satellite73 I thought the inverse of the function was f^-1(x)=-1n(7-x)

OpenStudy (anonymous):

how exactly do I find the domain, range, and asymptotes?